# Transparent and Coherent Procedural Mistake Detection

### v2

@@ 6.1 Multi-Step Reasoning in LMs & VLMs
LMs have exhibited impressive reasoning capabilities from prompting methods Wei et al. (2022); Kojima et al. (2022), later strengthened with multiple paths Wang et al. (2023b); Snell et al. (2024), tree-search Yao et al. (2023); Hao et al. (2023); Putta et al. (2024); Tian et al. (2024); Chen et al. (2024a); Zhang et al. (2024); Qi et al. (2024), and fine-tuning on reasoning chains from stronger LMs Wang et al. (2024); Gou et al. (2024); Muennighoff et al. (2025) and/or self-generated during reinforcement learning OpenAI et al. (2024a); DeepSeek-AI et al. (2025). Related to PMD, some work has investigated LMs’ reasoning about dependencies between physical procedures Bellos et al. (2024); Lal et al. (2024). Meanwhile, other work has applied iterative self-questioning approaches to deepen inquiry in other domains, e.g., medicine and fact verification Cattan et al. (2024); Li et al. (2025); Vladika et al. (2025).
In light of challenges in visual reasoning in VLMs Dai et al. (2023b); Li et al. (2023b); Guan et al. (2024), some work has proposed training-free strategies and training paradigms to reduce visual hallucination Wan et al. (2024); Leng et al. (2024); An et al. (2024); Ganz et al. (2024), and utilized LMs and other tools to generate intermediate questions and coordinate visual reasoning step by step You et al. (2023); Srinivasan et al. (2024); Chen et al. (2024b); Zhou et al. (2024); Zong et al. (2024); Cheng et al. (2024); Zhu et al. (2024); Sun et al. (2024); Jing and Rahman (2025). The latter stems from prior efforts in visual dialog Das et al. (2017); Kottur et al. (2019), which has been used to study and improve machine performance in various problem areas Chattopadhyay et al. (2017); de Vries et al. (2017); Uehara et al. (2022); Wang et al. (2022). Other works have assessed (V)LMs’ reasoning about procedures based on visual information Yang et al. (2018); Hendricks and Nematzadeh (2021); Jin et al. (2022); Yuksekgonul et al. (2023); Nguyen et al. (2024, 2025). In this work, we evaluated self-generated visual dialogs from VLMs in a novel problem of coherent PMD.
@@ 6.2 Leveraging NLI in Other NLP Tasks
NLI requires determining whether a premise text entails a hypothesis text, a reasoning challenge long studied in NLP Dagan et al. (2005). Many human-annotated resources have been compiled for NLI, thus significant progress has occurred in developing NLI models Storks et al. (2019). Prior work has leveraged this progress by using NLI models to improve the competence, confidence, and coherence of LMs in other tasks, such as dialog systems Dziri et al. (2019); Welleck et al. (2019), summarization Roit et al. (2023), VQA Srinivasan et al. (2024), and image captioning Cascante-Bonilla et al. (2024). Building on this, we adopt two NLI-based metrics to strengthen VLMs in the novel problem of coherent PMD.
@@ 7 Conclusion
In this work, we evaluated foundational VLMs on a challenging problem of coherent PMD, where visual questions and answers must be generated to drive success decisions. To evaluate these rationales, we leveraged an NLI model to define two coherence metrics, using them to encourage coherent question selection and generation through common interventions. Our results showed that VLMs do not generate coherent rationales off-the-shelf, but these interventions improve their coherence, with the former also improving accuracy, and the latter improving the efficiency of generating and extracting information from rationales, albeit creating tradeoffs between these aspects. Further, patterns in accuracy and coherence metrics revealed detailed performance insights. Ultimately, when choosing an approach for coherent PMD, accuracy, coherence, and efficiency must be weighed for the setting. For example, while online applications may require shorter dialog (even at a cost of accuracy), high-risk applications may prioritize accuracy and confidence. This work lays a foundation for future research in the application of VLMs to PMD and task guidance, and investigation of how these system goals interact.
8 Limitations
@@
One limitation of our coherent PMD problem formulation is the requirement of generating several pieces of information autoregressively, which would take several seconds in practical settings. This is not ideal for a problem like interactive task guidance, where responsiveness and the ability to intervene quickly to correct mistakes are important. However, rather than being applied frame by frame (which would likely not be feasible), we expect this process to be applied once at the end of procedure execution to verify the state of the environment, e.g., when the user asks a task guidance system to inform them of the next step of a recipe. Based on the results of our study, one could explore streamlined and specialized approaches to apply VLMs to a stream of video frames in a live online setting. For example, in preliminary experiments, we tried to generate questions once with a VLM, then answer them over a series of video frames, but we found this approach limited by the inability to adapt questions to previously gathered information, and the challenge of aggregating noisy VLM responses across time. We leave further investigation of such approaches for future work.
@@
Next, single video frames are limited in representing actions, which involve movement and state change. Our decision to focus on individual frames stemmed from preliminary experiments we performed with existing open-source VLMs for video understanding Lin et al. (2024); Li et al. (2023a), which are still in early stages. There, we found that they often confused information from frames in different regions of the video, preventing them from judging the final states of objects and reconciling this information with success of procedures, and resulting in poor performance. As such, our choice to focus on single images simplified the problem for current VLMs, enabling our experiments to begin building a meaningful understanding of their capabilities in PMD. To minimize the dependence on multiple frames in detecting mistakes, we applied several careful preprocessing steps to Ego4D-PMD (as discussed thoroughly in Section 2.3 and Appendix A). It is also worth noting that by not incorporating modalities beyond text and images, e.g., audio, the VLMs we studied are inherently limited in their capturing of physical information Yu et al. (2022); Zong et al. (2024). As the ability to reason over sequences of frames and other modalities evolves in state-of-the-art VLMs, future work can revisit this formulation and explore new approaches for them to reason over dynamic scenes.
@@ Acknowledgments
This work was supported in part by the DARPA PTG program HR00112220003. We would like to thank Megan Su, Ruixuan Deng, Fengyuan Hu, Andy Chung, Lu Wang, Richard L. Lewis, and the entire MSRP team for their helpful discussions, feedback, and insights.
@@ Appendix A Ego4D-PMD Data Curation Details
One natural type of mistake a user could make is not finishing a procedure. In addition to postcondition times, each video clip is annotated with a precondition time. Following a similar approach in Du et al. (2023), we can generate a mistake example by sampling a frame at the precondition time and pairing it with the video clip’s narration text. We expect that by doing this, the sampled frame will show the procedure at an incomplete state, and contain most of the same objects as the success example for the same clip. This poses a difficult challenge of identifying the key physical properties of the scene that would indicate completion.
@@
Mistakes also happen when a user applies the wrong type of action to an object, causing an unexpected state, as well as when a user uses the wrong object or ingredient in a procedure. Following this intuition, we generate additional easier mistake examples from each clip by matching each clip with other clips that have a mismatched verb, noun, or both. While each clip is annotated with verb and noun categories, these categories are coarse-grained, making it impossible to guarantee that two clips with the same verb or noun label actually involve the same verb or noun, thus preventing sampling clips that share the same verb or noun. Instead, we apply the AllenNLP [^19] https://allenai.org/allennlp semantic role labeler to each narration text to identify the key participants in each procedure. For each clip, we then attempt to sample the postcondition frames from three mismatched clips: one with a mismatched verb (but matching nouns), one with a mismatched noun (but matching verb), and one with a mismatched verb and noun. We then pair these frames with the source clip’s narration text, creating mistake examples with varying levels of overlap with the source clip. While it is not always possible to find every such alternative clip for each clip in Ego4D, we can usually find at least one of them.
@@
We perform several additional steps to ensure high-quality mistake detection task instances. First, we remove clips where the precondition and postcondition frames are overly similar (i.e., at least 0.95 cosine similarity). We remove clips that are too dark (i.e., where the mean of all normalized RGB values is less than 0.2). When sampling frames from source clips, we sample several candidates within a small range around the precondition or postcondition timestamp, then select the least blurry candidate by the variance of the images’ Laplacian. Some videos in Ego4D show the same action being performed over and over (e.g., “Roll a ball of dough”), which can make it difficult to determine whether the state of the environment shown in a clip is the result of the current procedure or a prior one (given only a single frame). While future work applying video-optimized VLMs for coherent PMD in long-horizon tasks will need to address this challenge, this adds an unnecessary complexity to an already challenging task for current VLMs. As such, we remove any clips such that the same procedure in the clip has already been performed previously in the video. Lastly, we remove a handful of videos in Ego4D that we notice to be corrupted or significantly distorted.
@@ Appendix B Rephrasing Questions and Answers for Coherence Evaluation
Appendix C Coherence Metrics Human Study Details
In this appendix, we collect human judgments for the relevance and informativeness metrics defined in Section 3. To achieve this, we randomly sampled the outputs for 50 iterations of LLaVA’s self-dialog from two combinations of approaches:
@@ C.1 Annotation Instructions
  • 4 (slightly informative/sure)
  • 5 (very informative/sure)
@@ C.2 Human Study Results
We presented each annotator with 10 randomly selected VLM-generated questions $Q^{\prime}$ and answers $A^{\prime}$, along with previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$. Annotators were instructed to rate the relevance (i.e., based on all the information we have, how sure is the annotator about whether the procedure was successfully completed) on a scale from 1-5 (least to most informative). Between the resulting 50 annotations and corresponding automated metrics, we found a weaker Spearman correlation Spearman (1904) of $\rho=0.33$ ($p=0.020$). Interestingly, if we multiply the automated informativeness metrics by the relevance for $Q^{\prime}$, $\mathcal{Q}$, and $\mathcal{A}$, we find a stronger Spearman correlation of $\rho=0.50$ ($p=0.00022$). This suggests that while informativeness does have a relationship with human judgments, when multiplying it by relevance this relationship is stronger and more significant. This might be because the concepts of relevance and informativeness are themselves related. Intuitively, in most cases, a relevant question should be informative, and an irrelevant question should be uninformative. Meanwhile, fine-tuned NLI models could theoretically score an answer to an irrelevant question as informative, and vice versa. When proposing a coherence-based question re-ranking strategy in Section 4.1.1, we incorporated the inductive bias reflected in these human judgments by multiplying informativeness by relevance. Later, in Section 5.3, we visualized the full distribution of relevance and informativeness on our evaluation data to better understand this issue.

### v3

@@ 6.1 Multi-Step Reasoning in LMs & VLMs
LMs have exhibited impressive reasoning capabilities from prompting methods Wei et al. (2022); Kojima et al. (2022), later strengthened with multiple paths Wang et al. (2023b); Snell et al. (2024), tree-search Yao et al. (2023); Hao et al. (2023); Putta et al. (2024); Tian et al. (2024); Chen et al. (2024a); Zhang et al. (2024); Qi et al. (2024), and fine-tuning on reasoning chains from stronger LMs Wang et al. (2024); Gou et al. (2024); Muennighoff et al. (2025) and/or self-generated during reinforcement learning OpenAI et al. (2024b); DeepSeek-AI et al. (2025). Related to PMD, some work has investigated LMs’ reasoning about dependencies between physical procedures Bellos et al. (2024); Lal et al. (2024). Meanwhile, other work has applied iterative self-questioning approaches to deepen inquiry in other domains, e.g., medicine and fact verification Cattan et al. (2024); Li et al. (2025); Vladika et al. (2025).
In light of challenges in visual reasoning in VLMs Dai et al. (2023b); Li et al. (2023b); Guan et al. (2024), some work has proposed training-free strategies and training paradigms to reduce visual hallucination Wan et al. (2024); Leng et al. (2024); An et al. (2024); Ganz et al. (2024), and utilized LMs and other tools to generate intermediate questions and coordinate visual reasoning step by step You et al. (2023); Srinivasan et al. (2024); Chen et al. (2024b); Zhou et al. (2024); Zong et al. (2024); Cheng et al. (2024); Zhu et al. (2024); Sun et al. (2024); Jing and Rahman (2025); Cheng et al. (2025). The latter follows prior work in visual dialog Das et al. (2017); Kottur et al. (2019), which has been used to study and improve machine performance in various problem areas Chattopadhyay et al. (2017); de Vries et al. (2017); Uehara et al. (2022); Wang et al. (2022). Other works have assessed (V)LMs’ reasoning about procedures based on visual information Yang et al. (2018); Hendricks and Nematzadeh (2021); Jin et al. (2022); Yuksekgonul et al. (2023); Nguyen et al. (2024, 2025). In this work, we evaluated self-generated visual dialogs from VLMs in a novel problem of coherent PMD.
@@ 6.2 Leveraging NLI in Other NLP Tasks
NLI requires judging whether a premise text entails a hypothesis text, a reasoning challenge long studied in NLP Dagan et al. (2005). Many human-annotated NLI resources have been created, thus significant progress has occurred in NLI Storks et al. (2019). Consequently, prior work has used NLI models to improve the competence, confidence, and coherence of LMs in other tasks, e.g., dialog systems Dziri et al. (2019); Welleck et al. (2019), summarization Roit et al. (2023), VQA Srinivasan et al. (2024), and image captioning Cascante-Bonilla et al. (2024). We similarly adopt two NLI-based metrics to strengthen VLMs in a new problem of coherent PMD.
@@ 7 Conclusion
In this work, we evaluated foundational VLMs on a challenging problem of coherent PMD, where visual questions and answers must be generated to drive success decisions. To evaluate these rationales, we leveraged an NLI model to define two coherence metrics, using them to encourage coherent question selection and generation through common interventions. Our results showed that VLMs do not generate coherent rationales off-the-shelf, but these interventions improve their coherence, with the former also improving accuracy, and the latter improving the efficiency of generating and extracting information from rationales, albeit creating tradeoffs between these aspects. Further, patterns in accuracy and coherence metrics revealed detailed performance insights, e.g., visual processing errors like object hallucination. Ultimately, when choosing an approach for coherent PMD, accuracy, coherence, and efficiency must be weighed for the setting. For example, while online applications may require shorter dialog (even at a cost of accuracy), high-risk applications may prioritize accuracy and confidence. This work lays a foundation for future research in the application of VLMs to PMD and task guidance, and investigation of how these system goals interact. Our code and data are available at https://github.com/sled-group/Transparent-Coherent-PMD.
Limitations
@@
One limitation of our coherent PMD problem formulation is the requirement of generating several pieces of information autoregressively, which would take several seconds in practical settings (as shown in Appendix E.3). This is not ideal for a problem like interactive task guidance, where responsiveness and the ability to intervene quickly to correct mistakes are important. However, rather than being applied frame by frame (which would likely not be feasible), we expect this process to be applied once at the end of procedure execution to verify the state of the environment, e.g., when the user asks a task guidance system to inform them of the next step of a recipe. Based on the results of our study, one could explore streamlined and specialized approaches to apply VLMs to a stream of video frames in a live online setting. For example, in preliminary experiments, we tried to generate questions once with a VLM, then answer them over a series of video frames, but we found this approach limited by the inability to adapt questions to previously gathered information, and the challenge of aggregating noisy VLM responses across time. We leave further investigation of such approaches for future work.
@@
Next, single video frames are limited in representing actions, which involve movement and state change. Our decision to focus on individual frames stemmed from preliminary experiments we performed with existing open-source VLMs for video understanding Lin et al. (2024); Li et al. (2023a), which are still in early stages. There, we found that they often confused information from frames in different segments of the video, preventing them from judging the final states of objects and reconciling this information with success of procedures, thus resulting in poor performance. As such, our choice to focus on single images simplified the problem for current VLMs, enabling our experiments to begin building a meaningful understanding of their capabilities in PMD. To minimize the dependence on multiple frames in detecting mistakes, we applied several careful preprocessing steps to Ego4D-PMD (as discussed thoroughly in Section 2.3 and Appendix A). It is also worth noting that by not incorporating modalities beyond text and images, e.g., audio, the VLMs we studied are inherently limited in their capturing of physical information Yu et al. (2022); Zong et al. (2024).
As the ability to reason over sequences of frames and other modalities evolves in state-of-the-art VLMs, future work can revisit this formulation and explore new approaches for them to reason over dynamic scenes. Specifically, we imagine that this task will look somewhat similar as VLMs mature. Given a VLM that can reliably extract information from multiple frames and from audio (e.g., about fine-grained motion, temperatures, and more) and express it in text, coherence metrics like those we explored in this work can still apply. In order to track this information reliably, a neuro-symbolic architecture may be required, e.g., a dynamic scene graph continuously updated across frames based on observations from a VLM and/or other neural vision models. Of course, there will always be some information that a VLM cannot extract from the environment even through vision and audio. For example, confirming amounts or weights of ingredients in a cooking setting may be theoretically possible by reading measurement tools as the human uses them, but this would be a highly difficult challenge. In such cases, a useful capability for a PMD system would be to ask the human user some questions (e.g., “Are you sure that was 3 tablespoons? That looked like a teaspoon.”), then incorporate the user’s answers into PMD and rationalization. Such a capability becomes a rich research problem of its own, as it would be important for the system not to ask too many questions Bao et al. (2023), requiring it to be able to identify and prioritize the cases where this is necessary, and to ask questions concisely and in a timely manner.
@@ Acknowledgments
This work was supported in part by the DARPA PTG program HR00112220003, NSF IIS-1949634, ARPA-H PARADIGM program 1AY2AX000062, the Microsoft Accelerate Foundation Models Research (AFMR) grant program, and computational resources and services provided by Advanced Research Computing at the University of Michigan, Ann Arbor. We would like to thank our anonymous reviewers, as well as Megan Su, Ruixuan Deng, Fengyuan Hu, Andy Chung, Lu Wang, Richard L. Lewis, and the entire MSRP team, for their helpful discussions, feedback, and insights. ChatGPT [^20] https://chatgpt.com/ was used for minor writing suggestions and tedious coding tasks (e.g., improving generated graphs). This research was funded, in part, by the U.S. Government. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government.
@@ Appendix A Ego4D-PMD Data Curation Details
One natural type of mistake a user could make is not finishing a procedure. In addition to postcondition times, each video clip is annotated with a precondition time. Following a similar approach to Du et al. (2023), we can generate a mistake example by sampling a frame at the precondition time and pairing it with the video clip’s narration text. We expect that by doing this, the sampled frame will show the procedure at an incomplete state, and contain most of the same objects as the success example for the same clip. This poses a difficult challenge of identifying the key physical properties of the scene that would indicate completion.
@@
Mistakes also happen when a user applies the wrong type of action to an object, causing an unexpected state, as well as when a user uses the wrong object or ingredient in a procedure. Following this intuition, we generate additional easier mistake examples from each clip by matching each clip with other clips that have a mismatched verb, noun, or both. While each clip is annotated with verb and noun categories, these categories are coarse-grained, making it impossible to guarantee that two clips with the same verb or noun label actually involve the same verb or noun, thus preventing sampling clips that share the same verb or noun. Instead, we apply the AllenNLP [^21] https://allenai.org/allennlp semantic role labeler to each narration text to identify the key participants in each procedure. For each clip, we then attempt to sample the postcondition frames from three mismatched clips: one with a mismatched verb (but matching nouns), one with a mismatched noun (but matching verb), and one with a mismatched verb and noun. To avoid incorrect mismatches, we omitted some verbs from this matching that are often prerequisite for other object interactions and thus likely to occur even in sampled videos thought to have mismatched verbs, such as movement verbs like “take” and “put.” We also verified that the verb-noun pair from the source video did not happen earlier in the retrieved video, ensuring that the correct physical states cannot be observed in the retrieved video from earlier actions. We then pair sampled frames with the source clip’s narration text, creating mistake examples with varying levels of overlap with the source clip. While it is not always possible to find every such alternative clip for each clip in Ego4D, we can usually find at least one of them.
@@
We perform several additional steps to ensure high-quality mistake detection task instances. First, we remove clips where the precondition and postcondition frames are overly similar (i.e., at least 0.95 cosine similarity). We remove clips that are too dark (i.e., where the mean of all normalized RGB values is less than 0.2). When sampling frames from source clips, we sample several candidates within a small range around the precondition or postcondition timestamp, then select the least blurry candidate by the variance of the images’ Laplacian. Some videos in Ego4D show the same action being performed repeatedly (e.g., “Roll a ball of dough”), which can make it difficult to determine whether the state of the environment shown in a clip is the result of the current procedure or a prior one (given only a single frame). While future work applying video-optimized VLMs for coherent PMD in long-horizon tasks will need to address this challenge, this adds an unnecessary complexity to an already challenging task for current VLMs. As such, we remove any clips such that the same procedure in the clip has already been performed previously in the video. Lastly, we remove a small number of videos in Ego4D that we notice to be corrupted or significantly distorted.
@@ Appendix B Rephrasing Questions and Answers for Coherence Evaluation
Appendix C Human Annotation Details
In this appendix, we collect human judgments for the relevance and informativeness metrics defined in Section 3, as well as PMD decisions.
C.1 Validating Coherence Metrics
To achieve this, we randomly sampled the outputs for 50 iterations of LLaVA’s self-dialog from two combinations of approaches:
@@ C.1.1 Annotation Instructions
You can also choose to mark "Instructions Unclear", which means that the sentence itself is not clear, so you’re not sure how to determine whether the procedure is successful. This should only be used in rare cases.
Some tips:
  • Only judge the relevance of the potential next question, not the previous questions (which may or may not be relevant).
  • A question may seem relevant to the task at hand, but you should consider it irrelevant if it can’t provide essential information to judge whether the task was completed successfully.
  • If a seemingly relevant question is redundant with previous questions, you may consider it less relevant.
  • Assume that the answer to the question won’t contradict the information you have from previous questions and answers. If previous questions and answers already contradict each other, consider whether this question could sway you one way or another.
  • The instructional text and questions may refer to "someone" or "a person"; always assume this is referring to yourself (the person performing the task).
  • The questions may refer to a "photo" or "image"; always assume this is referring to the video feed your friend would see through the video call.
@@
You can also choose to mark "Instructions Unclear", which means that the sentence itself is not clear, so you’re not sure how to determine whether the procedure is successful. This should only be used in rare cases.
Some tips:
  • Your task is to rate how sure you are, NOT whether you believe the procedure is successfully completed or not.
  • Consider all questions and answers as a whole; if you have contradictory information, this may reduce your sureness.
  • The instructional text and questions may refer to "someone" or a "person"; always assume this is referring to yourself (the person performing the task).
  • The questions may refer to a "photo" or "image"; always assume this is referring to the video feed your friend would see through the video call.
@@
  • 4 (slightly informative)
  • 5 (very informative)
@@ C.1.2 Annotation Results
We presented each annotator with 10 randomly selected VLM-generated questions $Q^{\prime}$ and answers $A^{\prime}$, along with previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$. Annotators were instructed to rate the relevance (i.e., based on all the information we have, how sure is the annotator about whether the procedure was successfully completed) on a scale from 1-5 (least to most informative). Between the resulting 50 annotations and corresponding automated metrics, we found a weaker Spearman correlation Spearman (1904) of $\rho=0.33$ ($p=0.020$). Interestingly, if we multiply the automated informativeness metrics by the relevance for $^{\prime}, 𝒬\mathcal{Q}$, and $\mathcal{A}$, we find a stronger Spearman correlation of $\rho=0.50$ ($p=0.00022$). This suggests that while informativeness does have a relationship with human judgments, when multiplying it by relevance this relationship is stronger and more significant. This might be because the concepts of relevance and informativeness are themselves related. Intuitively, in most cases, a relevant question should be informative, and an irrelevant question should be uninformative. Meanwhile, fine-tuned NLI models could theoretically score an answer to an irrelevant question as informative, and vice versa. When proposing a coherence-based question re-ranking strategy in Section 4.1.1, we incorporated the inductive bias reflected in these human judgments by multiplying informativeness by relevance. Later, in Section 5.3, we visualized the full distribution of relevance and informativeness on our evaluation data to better understand this issue.
C.2 Human PMD Accuracy
To create a reference point for PMD accuracy measurements of VLMs and a proxy for the Ego4D-PMD data quality and objectiveness, we performed a small human annotation study for the task of PMD. Specifically, we randomly selected 100 examples from the testing data, and asked 3 annotators to judge whether each example’s frame showed a successful execution of the actions described by its procedural text. We recruited a total of 12 annotators, each of which labeled 25 examples. All annotators were proficient English speakers with conferred or in-progress undergraduate degrees, and again were uncompensated authors of the work (not the first author) and/or peers of the authors. Annotators were given the following instructions:
You will be shown 25 pairs of task instructions and egocentric (POV) photos from people’s perspective. Based on each photo, you will be asked whether the person has successfully completed the given task. You might see a range of situations, including the person following the instructions perfectly, making a minor mistake, or doing something completely different.
Some guidelines:
  • Based on the task instructions, you should look for one or more objects that you expect to be involved in the task. Based on what they look like and where they’re located, you should judge whether the task has been successfully completed.
    • For example, if the task is to "slice an apple," and you only see whole unsliced apples in the photo, it should be labeled a failure ("No - They did something wrong").
    • Or if you do see a sliced apple, it should be labeled a success ("Yes - Task completed successfully").
  • It’s possible the photo doesn’t provide enough information to decide if the task is complete. If you don’t see any objects required for the task, only see part of an object, or the photo is blurry/low quality, it may or may not still be a success. Make your best guess based on what you do see in the image.
    • You will also have a chance to indicate your confidence.
For each example, annotators were shown the example’s video frame and procedural text, then asked:
Did the person successfully complete the task?
  • No - They did something wrong
  • Yes - Task completed successfully
They were also asked to indicate their confidence on a scale from 1 to 5:
How confident are you in your answer?
  • 1 - Not confident at all
  • 2 - Slightly confident
  • 3 - Moderately confident
  • 4 - Very confident
  • 5 - Extremely confident
Pairs of questions for examples were presented in a random order for each annotator. To calculate human accuracy, we explored two approaches: (1) calculating the accuracy over the 3 annotators for each example then averaging over all examples, and (2) taking the majority judgement from the 3 annotators then calculating the accuracy based on those. The mean human accuracy was 71.3%, and the majority human accuracy was 72.0%, compared to the best model result of 67.8% in the coherent PMD setting. [^23] GPT-4o achieves 69.2% accuracy in the rationale-free evaluation presented in Appendix E.6, which is closer to human performance. However, it is important to note that this evaluation excludes the step of forming a rationale, which we assume humans do implicitly in making their decisions. This shows that VLMs still lag slightly behind humans in accurately detecting mistakes, and if human performance is to be considered an upper bound, coherently rationalizing decisions will remain a crucial step in this problem. Meanwhile, the mean human confidence rating was 3.52 out of 5, suggesting humans were usually moderately to very confident in their decisions. Lastly, the mean Cohen’s $\kappa$ Cohen (1960) for inter-annotator agreement was 0.572, suggesting moderate agreement on decisions.